# RealLiFe: Real-Time Light Field Reconstruction via Hierarchical Sparse Gradient Descent

### v3

@@ 1 Introduction
With the rapid development of XR / naked eye 3D display devices, there is an increasing demand for live light field video rendering techniques that can record and stream light field video in real time, allowing viewers to interactively change the viewpoint and focus of the video. Even with the latest breakthroughs in light field reconstruction[1, 2] and neural radiance fields[3, 4], it remains a challenging task.
@@
Figure 2: The derivation of MPI sparse gradients and the optimization process. (a) An MPI plane $M_{d}$ at depth $d$ can be decomposed into an RGB plane $c_{d}$ and an alpha (transparency) plane $\alpha_{d}$. (b) The multi-plane alpha gradients $\mathcal{A}_{d}$ are computed from multi-plane alpha planes by calculating the gradient of $M$ corresponding to $c_{d}$, and $\mathcal{A}_{d}$ is then sparsified by discarding small gradients, which is used to refine the MPI using our proposed sparse gradient descent strategy.
@@
  • Methods relying on implicit scene representation, such as NeRF [3] and its derivatives [10, 11, 12, 13] can produce view synthesis results with high visual quality, but require long-time per-scene optimization and dense view inputs. While some generalization techniques [14, 15, 16, 5, 6, 17] can generate novel views that are generalizable with the help of multi-view stereo (MVS) geometry rules, it is still not fast enough for them to achieve real-time performance due to high memory usage and complex computation of intermediate cost volumes.
@@
  • Methods that rely on explicit scene representation, such as Plenoxels[18], PlenOctrees[19] and DVGO[20] enclose a 3D scene in a voxel grid and optimize the radiance fields within it. While such methods can reproduce intricate geometric and texture details at high grid resolutions, they can not generalize to new scenes without extensive per-scene optimization. Although some methods with learned features [21, 9, 22, 1] can generalize well to unseen scenes from sparse views, they still require a significant amount of time to infer a scene due to the heavy computational burden of processing large multi-plane image (MPI) and plane sweep volume (PSV).
@@
  • Methods relying on surrogate geometry are computationally efficient by sampling only around the estimated geometry scaffolds, such as depth maps [23] and surface mesh. However, these methods may fail if the precomputed geometry scaffolds are incorrectly estimated for semi-transparent surfaces or thin plates; therefore, they are not very robust for complex scenarios.
Based on the above observations, we conclude that live light field reconstruction remains a challenging task for the following reasons:
1) Recovering a 3D light field from 2D images is a difficult ill-posed inverse problem that necessitates either dense input views or geometric priors learned from data;
2) Implicit representations save memory but are computationally complex, resulting in a long inference time. Explicit representations offer faster rendering speed but require significant memory and time for constructing the light field;
3) High-quality light field reconstruction necessitates sophisticated optimization to handle intricate geometry details such as semi-transparent surfaces and thin structures.
To overcome the challenges, we propose RealLiFe, a novel method that employs Hierarchical Sparse Gradient Descent (HSGD) to produce high-quality light fields in real time from sparsely sampled views.
Unlike previous methods[21, 9, 1] that directly process and generate a full MPI, our method adopts a hierarchical approach, progressively refining it from low resolution to high resolution. This approach draws inspiration from the principles of Multi-View Stereo (MVS), where intricate geometric details can be iteratively enhanced with high-resolution features and well-established low-resolution geometric structures[24, 25, 26, 27, 28, 29]. We extend this concept from geometry to appearance, claiming and proving that the visual details within a light field can also be iteratively improved, given access to high-resolution captures and well-structured low-resolution representations.
[…]
Before the design for efficiency, it is imperative to address the fundamental optimization process for deriving an MPI from sparse views. This task poses a significant challenge, as it involves solving an inverse problem. Conventionally, this is accomplished by iteratively minimizing the discrepancies between the predicted and observed measurements using analytic gradient descent, which can be a time-consuming process. However, DeepView[1] introduces a novel approach by incorporating the learned gradient descent strategy[30, 31]. This strategy enables the rapid generation of an MPI in only a few iterations by explicitly computing MPI gradients and relying on the neural network to perform adaptive gradient descent. We also apply this strategy for rapid gradient descent, but, in contrast to DeepView [1], which utilizes full-sized and full-channel MPI gradients, our approach only employs sparsely gradients that align with the scene contents. This choice arises from the observation that gradients within vacuous regions either insignificantly affect or negligibly contribute to the final rendering quality.
Based on our observations, we present the fundamental stages of the proposed Hierarchical Sparse Gradient Descent (HSGD), as depicted in Fig. 2. For example, in one stage, the multi-plane gradients $\mathcal{A}_{d}$ are initially obtained by computing the gradients of the MPI with respect to its RGB channel, as outlined in DeepView[1]. Subsequently, $\mathcal{A}_{d}$ undergoes sparsification by removing small-value gradients. Finally, the sparse gradient descent module updates the MPI based on the sparse gradients.
@@
Figure 4: The overview of RealLiFe. First, the PSV is constructed using multiview images by homographic warping, and the PSV is then downsampled hierarchically at multiple resolutions. The output MPI is then generated in several iterations. Initially, the lowest-resolution PSV is fed to a 3D CNN to extract a coarse-level MPI. The PSV and the upsampled MPI both go through the Sparse Gradient Descent module for a refined higher-resolution MPI. This loop is repeated until the resolution of the output MPI reaches the same resolution as that of the input images. Finally, a novel view can be easily rendered from the MPI using the repeated over operator[33].
@@ 3.2 Initial MPI Generation
Hierarchical PSV Construction. Given $N$ source view images ${I}_{i=1}^{N}$ of size $H\times W\times 3$ and their corresponding camera poses $[K_{i},R_{i}|t_{i}]$, as well as for the reference view $I_{r}$, we construct the plane sweep volume $P$ leveraging inverse homographic warping:
@@ 3.3 Hierarchical Sparse Gradient Descent
Hierarchical sparse gradient descent is a coarse-to-fine approach that progressively refines the MPI using sparse gradients. This design is supported by two key insights: Coarse-to-fine Appearance Refinement and Sparse Gradients Suffice for Light Field Reconstruction. The coarse-to-fine geometry refinement has previously demonstrated its efficacy in various MVS works[24, 25, 26, 27, 28, 29]. In this context, a low-resolution geometric scaffold can be rapidly derived and employed as a prior to guide the subsequent high-resolution detail refinement process. Similarly, the task of deriving an MPI involves allocating scene contents to their respective depth planes. So the hierarchical pipeline is well-suited for MPI generation, offering a structured foundation upon which the learned gradient descent algorithm can iteratively refine appearance details. And the choice of sparse gradients stems from the empirical observation in Fig. 6 that after only one iteration of network forwarding, above half of the voxels in an MPI become empty. Hence, any subsequent optimization targeting these empty regions does not lead to appreciable enhancements in rendering quality but instead incurs unnecessary computational overhead.
Figure 6: Proportions of empty pixels of individual MPI planes throughout optimization iterations. (a) A rendered image of RealLife (without using HSGD) with 40 MPI planes. (b) A subset of MPI planes (i.e., the 12th, 16th, and 20th MPI planes) of three iterations. (c) Proportions of empty pixels in individual MPI planes. (pixels characterized by $\alpha<0.1$ are considered empty, as their contribution to rendering quality is marginal.)
@@
where $\widetilde{P}$ is the PSV constructed with rendered images at each source view from the generated MPI, and $\mathcal{T}^{w^{-1}}_{i}$ is the warped MPI gradient respective to MPI’s $\alpha$ channel. But we simplify the gradients to be in the format of 10:
~$\widetilde{\nabla}=[P,\mathcal{A}^{w^{-1}}_{1},...,\mathcal{A}^{w^{-1}}_{N}],$ (10)
@@
To compute $\mathcal{A}^{w_{i}^{-1}}$ for a single source view ${i}$, the MPI of the reference view is initially warped to source view $i$ to obtain $M_{i}$. Then the alpha gradient $\mathcal{A}_{i_{d}}$ corresponding to depth $d$ is calculated by the following equation:
@@
Gradient Sparsification. To safely sparsify gradient components without adversely affecting the final rendering quality, it becomes imperative to establish criteria for the exclusion of voxels that impart minimal contributions. A commonly adopted strategy in various Multiview Stereo (MVS) studies[24, 25, 26] is to reduce the number of depth layers in the full cost volume from $D$ to $k$, where $k$ is usually an odd number, composed of the plane with the highest probability and its adjacent $22\times(k-1)/$ 2 layers. However, this MVS-sampling strategy primarily relies on surface-centric geometric considerations, which might be too aggressive to capture some light field details like semi-transparent objects and thin plates. If we revisit Eq. 2, it can be rewritten as:
@@ 3.4 Training
$L_{r}=\sum_{i=1}^{l}\lambda_{i}(L_{VGG}(I_{r_{i}},\tilde{I}_{r_{i}})+\mu\sum_{j=1}^{N}L_{VGG}(I_{j_{i}},\tilde{I_{j_{i}}})),$ (15)
@@ 4 Implementation Details
The generalized model was trained using an RTX 3090 GPU with the Adam optimizer[47]. The initial learning rate was configured at $31e-$ 3, with a reduction by half if the training loss consistently decreased for a continuous span of 10 epochs. The model underwent training for 1000k iterations, taking approximately 13 hours to complete. The trainable components within the entire pipeline included the 3D CNNs responsible for converting the input plane sweep volume into a multi-plane image (as depicted in Fig. 4) and refining the sparse MPI gradient (as illustrated in Fig. 5). The structure of the 3D CNN, as shown in Fig. 8, comprised 5 basic convolutional blocks (a convolution layer followed by a batchnorm layer and a relu layer) and 1 1x1x1 3D convolution layer. To further improve the efficiency of our method, we optimize to run our pipeline with TensorRT [^1] https://developer.nvidia.com/tensorrt.
@@
Figure 8: The backbone 3D CNN of RealLiFe, which contains only three convolution layers.
@@
Category Model Generation Time, sec Offline Rendering, fps 3D Display Efficiency, fps SWORD Real Forward-Facing Shiny
@@
Configuration Soft3D[49] ${}^{+}$[21] DeepView[1] RealLiFe RealLiFe-2I
Small baseline 0.9260 0.8884 0.9541 0.9396 0.9286
Medium baseline 0.9300 0.8874 0.9544 0.9317 0.9159
Large baseline 0.9312 0.8673 0.9485 0.9180 0.9052
Generation Time, sec - - $\sim$20.00 0.044 0.037
@@ 5.1 Baselines, datasets and metrics
Baselines. In order to evaluate the visual performance and real-time efficiency of our method, we compare it with four generalizable novel view synthesis methods: IBRNet[5], MVSNeRF[6], and ENeRF[7]. Additionally, we compare our method with five light field reconstruction methods, namely Stereo Magnification (StereoMag)[21], Soft3D[49], MLI[8], and DeepView[1]. Among the baseline methods, IBRNet, MVSNeRF and ENeRF conduct extra finetuning experiments to improve performance. However, finetuning takes so much time that it is incompatible with real-time light field reconstruction, thus we only compare the results without finetuning. LLFF[9] is excluded as a baseline method because the number of input views for the released model is set fixed to 5. Additionally, LLFF fuses multiple MPIs to render novel views, leveraging information of up to 10-20 input views, making it a method with dense view inputs. However, it is essential to acknowledge that LLFF[9] remains a robust approach for generating high-quality light fields offline.
@@ 5.3 Quantitative Results
Compared with offline methods, our model RealLiFe and RealLiFe-2I stands out in terms of visual metrics, achieving top-1 or top-2 rankings on three datasets. And our default model RealLiFe strikes a balance between the rendering quality and efficiency. Our models, on average, achieve a 100×100$\times$100 × faster generation speed compared to other offline methods. Compared with the online method. Our models are a little bit faster than ENeRF[7] in generation time; however, our method exhibits superior visual metrics on three datasets. When considering the metric of 3D Display Efficiency, our method stands out as the only one capable of supporting real-time light field video display. This is due to its high temporal efficiency in both light field reconstruction and offline rendering speed.
Tab. II lists the quantitative results on the Spaces evaluation set, where we can see that DeepView generates the highest-quality rendering results overall, followed by our method and Soft3D[49]. The decrease in rendering quality may be attributed to the fact that our method depends on network connections to propagate visibility information. As we increase the baseline, there is a corresponding increase in the number of network connections required to effectively propagate visibility, as noted in [1]. However, in terms of time efficiency, our method is about 400 times faster than DeepView. The Generation Time of Soft3D and ${}^{+}$ are left as ”-” because Soft3D is not open-sourced and ${}^{+}$ is a modified version by the authors of DeepView, so that we can not evaluate their time efficiency accurately. To summarize, we have the following observations:
  • Compared with offline novel view synthesis methods[5, 6] and offline light field reconstruction methods[21, 49, 1, 8], our approach achieves around 100×100$\times$100 × speedup ratio with superior or comparable visual performance.
@@ 5.5 Ablation Study
Number of planes. When comparing the configurations of the default model, D=16 (fewer planes), and D=64 (fewer planes) in Tab. III, it becomes evident that employing too few planes results in a substantial deterioration in rendering quality, albeit compensating for expedited generation speed. Conversely, an increase in the number of planes yields only marginal enhancements in rendering quality while significantly diminishing temporal efficiency. This observation suggests the presence of a potential saturation point for the number of MPI planes. Beyond this point, the incremental addition of planes may not yield linear improvements in performance.
@@
Spasification criterion. In Section 3.3, we referenced an MVS[24, 25, 26, 27, 28, 29]-style gradient sparsification approach that involves selecting the layer with the highest alpha gradient along the depth axis as well as its adjacent $(k-1)/2$ layers to form a sparsified volume of $k$ layers of gradient data. This strategy, however, only takes geometry into consideration and may not adequately capture certain light field intricacies, such as semi-transparent objects and thin structures. An overall comparison in Table III reveals that the MVS-sampling strategy exhibits lower rendering quality when compared with the default model.
@@
$D=16$(fewer planes) 54.64 24.39 0.79 0.19
$D=64$(more planes) 28.57 25.48 0.84 0.15
$L=2$(2 iterartions) 45.45 25.33 0.84 0.16

### v4

@@ 1 Introduction
With recent advances in extended reality (XR) and autostereoscopic 3D display technologies, there is an urgent demand for real-time light field reconstruction systems capable of generating high-quality novel views from sparse input images. This technological imperative drives critical applications including low-latency 3D telepresence systems, interactive augmented reality interfaces, and immersive virtual reality platforms with adaptive viewpoint rendering capabilities.
Despite significant progress in neural light field reconstruction [1, 2] and neural radiance field representations [3, 4], achieving photorealistic rendering quality while maintaining real-time constraints for interactive applications remains challenging. Current approaches face fundamental trade-offs between visual performance and computational efficiency.
@@
  • Implicit representation methods, such as NeRF [3] and its derivatives [10, 11, 12, 13], produce high-quality view synthesis but require extensive per-scene optimization and dense view inputs. While generalization techniques [14, 15, 16, 5, 6, 17] can generate generalizable novel views using multi-view stereo geometry rules, they remain computationally intensive due to high memory usage and complex cost volume computations.
@@
  • Explicit representation methods, including Plenoxels[18], PlenOctrees[19], and DVGO[20], encapsulate 3D scenes in voxel grids and optimize radiance fields within them. While capable of reproducing intricate geometric and texture details at high resolutions, they cannot generalize to new scenes without extensive per-scene optimization. Methods with learned features [21, 9, 22, 1] achieve better generalization from sparse views but suffer from heavy computational burdens when processing large multi-plane images (MPI) and plane sweep volumes.
@@
  • Surrogate geometry methods improve computational efficiency by sampling around estimated geometry scaffolds, such as depth maps [23] and surface meshes. However, they may fail when geometry scaffolds are incorrectly estimated for semi-transparent surfaces or thin structures, limiting their robustness in complex scenarios.
These limitations reveal fundamental trade-offs in existing approaches: implicit methods demand high computational resources, explicit methods struggle with generalization, and surrogate geometry methods lack robustness without reliable initial geometric structures.
To address these challenges, we propose Hierarchical Sparse Gradient Descent (HSGD), which achieves a good trade-off between quality and efficiency in both training and inference stages. And our method leverages the potential geometric and color correspondences in the input plane sweep volumes to achieve generalization, which is free from using some explicit surrogate geometric structures.
Our key insight is that Multi-plane Images exhibit inherent sparsity in their gradient structure—most gradients have negligible magnitudes and contribute minimally to final rendering quality. By selectively processing only the most significant gradients, HSGD achieves substantial computational acceleration while maintaining visual fidelity. Specifically, The efficiency is achieved by the proposed 2-step processing. Firstly, the multi-plane gradients $\mathcal{A}_{d}$ are initially obtained by computing the gradients of the MPI with respect to its RGB channel, as outlined in DeepView[1]. Subsequently, $\mathcal{A}_{d}$ undergoes sparsification by removing small-value gradients. Finally, the sparse gradient descent module updates the MPI based on the sparse gradients.
@@
Figure 3: The overview of RealLiFe.(a) Initial MPI Generation: First, the PSV is constructed using multiview images by homographic warping, and the PSV is then downsampled hierarchically at multiple resolutions. The output MPI is then generated in several iterations. Initially, the lowest-resolution PSV is fed to a 3D CNN to extract a coarse-level MPI. (b) Hierarchical Sparse Gradient Descent: The PSV of multi-resolutions and the upsampled MPI both go through the Hierarchical Sparse Gradient Descent module for a refined higher-resolution MPI. Finally, a novel view can be easily rendered from the MPI using the repeated over operator[25].
@@ 3.2 Initial MPI Generation
Hierarchical PSV Construction. Given $N$ source view images ${I}_{i=1}^{N}$ of size $H\times W\times 3$ and their corresponding camera poses $K_{i},R_{i}|t_{i}]$, as well as for the reference view $I_{r}$, we construct the plane sweep volume $P$ leveraging inverse homographic warping:
@@ 3.3 Hierarchical Sparse Gradient Descent
Hierarchical sparse gradient descent is a coarse-to-fine approach that progressively refines the MPI using sparse gradients. This design is supported by two key insights: Coarse-to-fine Appearance Refinement and Sparse Gradients Suffice for Light Field Reconstruction. The coarse-to-fine geometry refinement has previously demonstrated its efficacy in various MVS works[40, 41, 42, 43, 44, 45]. In this context, a low-resolution geometric scaffold can be rapidly derived and employed as a prior to guide the subsequent high-resolution detail refinement process. Similarly, the task of deriving an MPI involves allocating scene contents to their respective depth planes. So the hierarchical pipeline is well-suited for MPI generation, offering a structured foundation upon which the learned gradient descent algorithm can iteratively refine appearance details. As analyzed in [46], the light field of a scene can be sufficiently represented by a single dominate plane given accurate depth observations, which indicate that the multi-plane image representation of light field is redundant. Experiments in Fig. LABEL:fig_sparseScene further demonstrates that after only one iteration of network forwarding, above half of the voxels in an MPI become empty. Hence, any subsequent optimization targeting these empty regions does not lead to appreciable enhancements in rendering quality but instead incurs unnecessary computational overhead.
@@
where $\widetilde{P}$ is the PSV constructed with rendered images at each source view from the generated MPI, and $\mathcal{T}^{w^{-1}}_{i}$ is the warped MPI gradient respective to MPI’s $\alpha$ channel. But we simplify the gradients to be in the format of Eq. 10:
$\widetilde{\nabla}=[P,\mathcal{A}^{w^{-1}}_{1},...,\mathcal{A}^{w^{-1}}_{N}],$ (10)
@@
To compute $\mathcal{A}^{w_{i}^{-1}}$ for a single source view i{i}, the MPI of the reference view is initially warped to source view $i$ to obtain $M_{i}$. Then the alpha gradient $\mathcal{A}_{i_{d}}$ corresponding to depth $d$ is calculated by the following equation:
@@
Gradient Sparsification. To safely sparsify gradient components without adversely affecting the final rendering quality, it becomes imperative to establish criteria for the exclusion of voxels that impart minimal contributions. A commonly adopted strategy in various Multiview Stereo (MVS) studies[40, 41, 42] is to reduce the number of depth layers in the full cost volume from $D$ to $k$, where $k$ is usually an odd number, composed of the plane with the highest probability and its adjacent $2\times(k-1)/2$ layers. However, this MVS-sampling strategy primarily relies on surface-centric geometric considerations, which might be too aggressive to capture some light field details like semi-transparent objects and thin plates. If we revisit Eq. 2, it can be rewritten as:
@@ 3.4 Training
$),L_{r}=\sum_{i=1}^{l}\lambda_{i}(L_{VGG}(I_{r_{i}},\tilde{I}_{r_{i}})+\mu\sum_{j=1}^{N}L_{VGG}(I_{j_{i}},\tilde{I_{j_{i}}}$)), (15)
@@ 4 Implementation Details
The generalized model was trained using an RTX 3090 GPU with the Adam optimizer[48]. The initial learning rate was configured at 1e31e-3, with a reduction by half if the training loss consistently decreased for a continuous span of 10 epochs. The model underwent training for 1000k iterations, taking approximately 13 hours to complete. The trainable components within the entire pipeline included the 3D CNNs responsible for converting the input plane sweep volume into a multi-plane image (as depicted in Fig. 3) and refining the sparse MPI gradient (as illustrated in Fig. 4). The structure of the 3D CNN, as shown in Fig. 6, comprised 5 basic convolutional blocks (a convolution layer followed by a batchnorm layer and a relu layer) and 1 1x1x1 3D convolution layer. To further improve the efficiency of our method, we optimize to run our pipeline with TensorRT [^1] https://developer.nvidia.com/tensorrt.
@@
Figure 6: The backbone 3D CNN of RealLiFe, which contains only six convolution layers.
@@
Category
Model
Generation
Time, sec
Offline
Rendering, fps
3D Display
Efficiency, fps
SWORD Real Forward-Facing Shiny
@@
Configuration Soft3D[50] StereoMag+[21] DeepView[1] DeepView(Re-Imp) RealLiFe RealLiFe-2I
Small baseline 0.9260 0.8884 0.9541 0.8912 0.9396 0.9286
Medium baseline 0.9300 0.8874 0.9544 0.8857 0.9317 0.9159
Large baseline 0.9312 0.8673 0.9485 0.8743 0.9180 0.9052
Generation Time, sec - - $\sim$20.00 $\sim$20.00 0.044 0.037
@@ 5.1 Baselines, datasets and metrics
Baselines. In order to evaluate the visual performance and real-time efficiency of our method, we compare it with four generalizable novel view synthesis methods: IBRNet[5], MVSNeRF[6], and ENeRF[7]. Additionally, we compare our method with four light field reconstruction methods, namely Stereo Magnification (StereoMag)[21], Soft3D[50], MLI[8], and DeepView[1]. Among the baseline methods, IBRNet, MVSNeRF and ENeRF conduct extra finetuning experiments to improve performance. However, finetuning takes so much time that it is incompatible with real-time light field reconstruction, thus we only compare the results without finetuning. LLFF[9] is excluded as a baseline method because the number of input views for the released model is set fixed to 5. Additionally, LLFF fuses multiple MPIs to render novel views, leveraging information of up to 10-20 input views, making it a method with dense view inputs. However, it is essential to acknowledge that LLFF[9] remains a robust approach for generating high-quality light fields offline.
@@ 5.3 Quantitative Results
Compared with offline methods, our model RealLiFe and RealLiFe-2I stands out in terms of visual metrics, achieving top-1 or top-2 rankings on three datasets. And our default model RealLiFe strikes a balance between the rendering quality and efficiency. Our models, on average, achieve a $100\times$ faster generation speed compared to other offline methods. Compared with the online method. Our models are a little bit faster than ENeRF[7] in generation time; however, our method exhibits superior visual metrics on three datasets. When considering the metric of 3D Display Efficiency, our method stands out as the only one capable of supporting real-time light field video display. This is due to its high temporal efficiency in both light field reconstruction and offline rendering speed.
Tab. II lists the quantitative results on the Spaces evaluation set, where we can see that DeepView generates the highest-quality rendering results overall, followed by our method and Soft3D[50]. The decrease in rendering quality may be attributed to the fact that our method depends on network connections to propagate visibility information. As we increase the baseline, there is a corresponding increase in the number of network connections required to effectively propagate visibility, as noted in [1]. We have also reported the results of our re-implemented version of DeepView, however, the results show some gaps with the original DeepView, this is mainly due to the visual discontinuities between small MPI patches. Since DeepView require high GPU memory to infer a full-sized MPI, so we have to infer small MPI patch by patch and then merge them all into a complete one, and this process leads to performance drop in our re-implemented version. However, in terms of time efficiency, our method is about 400 times faster than DeepView. The Generation Time of Soft3D and StereoMag+ are left as ”-” because Soft3D is not open-sourced and StereoMag+ is a modified version by the authors of DeepView, so that we can not evaluate their time efficiency accurately. To summarize, we have the following observations:
  • Compared with offline novel view synthesis methods[5, 6] and offline light field reconstruction methods[21, 50, 1, 8], our approach achieves around $100\times$ speedup ratio with superior or comparable visual performance.
@@ 5.5 Ablation Study
Number of planes. When comparing the configurations of the default model, D=16 (fewer planes), and D=64 (more planes) in Tab. III, it becomes evident that employing too few planes results in a substantial deterioration in rendering quality, albeit compensating for expedited generation speed. Conversely, an increase in the number of planes yields only marginal enhancements in rendering quality while significantly diminishing temporal efficiency. This observation suggests the presence of a potential saturation point for the number of MPI planes. Beyond this point, the incremental addition of planes may not yield linear improvements in performance.
@@
Spasification criterion. In Section 3.3, we referenced an MVS[40, 41, 42, 43, 44, 45]-style gradient sparsification approach that involves selecting the layer with the highest alpha gradient along the depth axis as well as its adjacent (k1)/2(k-1)/2 layers to form a sparsified volume of $k$ layers of gradient data. This strategy, however, only takes geometry into consideration and may not adequately capture certain light field intricacies, such as semi-transparent objects and thin structures. An overall comparison in Table III reveals that the MVS-sampling strategy exhibits lower rendering quality when compared with the default model.
HSGD full-sized infer v.s. DeepView patch-based infer. Leveraging our proposed Hierarchical Sparse Gradient Descent (HSGD) strategy, we are able to generate a full-sized MPI in a single forward pass. To examine the impact of this design, we also evaluated a patch-based inference configuration similar to DeepView: training and inferring on smaller patch-sized MPIs, cropping the edges of each patch, and merging them to form a complete full-sized MPI. For this comparison, we disabled the HSGD strategy, inferred four patches per full-sized MPI, and merged them into one. As shown in Table III, the “w/o HSGD” configuration exhibits a significant drop in efficiency, as it requires four forward passes and additional processing to crop and merge the patches. Moreover, these steps fail to completely eliminate visual discontinuities between patches, resulting in noticeable degradation in both efficiency and quality.
@@
D=16D=16(fewer planes) 54.64 24.39 0.79 0.19
D=64D=64(more planes) 28.57 25.48 0.84 0.15
L=2L=2(2 iterartions) 45.45 25.33 0.84 0.16